Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night Report 2014-09-15

Observers:  Morgan Fraser, Chris Ashall
TAT: Joe Anderson, Annalisa de Cia
Data reducers: Nic Walton, Nadia Blagorodnova

Conditions:
Thin to moderate clouds visible all directions at twilight
UT 22.41:    wind speed 10m/s, seeing 1-1.5", pointing restrictions to north-east



UT 22.59:    Twilight
                   B and R Sky Flats

UT 23.10:    Photometric standard EG274 

UT 23.33:    ASSAN-14gs classification, not a SN, Balmer lines in absorption at z=0. 1x300 gr13. 

UT 23.50:    SN 2009ip followip transparency very poor. Marginal detection in V and i, clear detection in R

UT 00.07:    LSQ14doz followp BVRI imaging, seen in all 4 bands

UT 00.16:    Transparency has improved, only thin clouds

UT 00.37:    No image analysis - TO set focus

UT 00.40:    LSQ14exl classification - acquisition image taken, target too faint to be seen, hence no spectrum was observed.

UT 00.52:    LSQ14fhd classification, SN Ia, pre-maximum, low signal to noise, z=0.06. gr13 1x900.

UT 01.27:    CSS140901-215624+190642 classification, low S/N (~5) spectrum is moderately blue and featurless. Unknown.

UT 01.56:    LSQ14eer followup, gr11 2x1800. Spectra had very low S/N - unsure if this is due to patchy clouds or target being fainter than expected, Took imaging later in night to check if SN is still visible.

UT 03.02:    LSQ14eez followip UBVRi imaging obtained

UT 04.20:    LSQ14eer followup - SN is still visible in BV imaging. 

UT 04.35:    SN 2009ip followup BVRi imaging repeated. Clouds not affecting imaging now, 09ip clearly deteceted in images.

UT 05:16:    LSQ14efd followup UBVRi photometry obtained

UT 05.48:    LSQ14efd followup 3x1800 gr13 spectra taken. 

UT 07.21:    ASASSN-14ha followup  1x1200 g11, 1x900 gr16

UT 07.21:    Transparency thin clouds visible

UT 08.16:    LSQ14fgo classification 1x900 gr13, blue and featureless spectrum

UT 08.32:    OGLE071 classification type II z=0.032

UT 08:56:    LSQ14fgo classification. Took 1x1800 gr11 spectrum to improve on initial classification spectrum. Still blue and featureless, although perhaps Balmer break in blue?

UT 09.23:     Twilight 

UT 09.27:    Spectrophotometric Standard LTT 7379 

UT 09.55:    CSS140914-010107-101840 classification spectra taken, no signal, it was saturated as the sky was too bright.

No sky flats taken due to moon illumination and bad transparency. 



            






Comments